Eddie is three times as old dave. seven years ago,eddie was five times old as dave.how old are they now?
Answer by Maths68(1474)

Let
x=Eddie’s current age
y=Dave’s current age
Eddie is three times as old Dave means
x=3y…………..(1)
Seven years ago means we have to subtract 7 from both ages
x-7=y-7
then
Eddie was five times old as Dave
x-7=5(y-7)
x-7=5y-35
x=5y-35+7
x=5y-28…………(2)
Put the value of x from (1) to (2)
x=5y-28
3y=5y-28
3y-5y=-28
-2y=-28
Divide by -2 both sides of above equation
-2y/-2=-28/-2
y=14
Put the value of y in equation (1)
x=3y
x=3(14)
x=42
Eddie is 42 and Dave is 14 years old.